OpenStudy (anonymous):

A girl has 8 coins in dimes and quarters. Their value is $1.55. How many quarters does she have?

OpenStudy (anonymous):

d+q=8 number of coins .10d+.25q=1.55 (value) 10d+25q=155 d=8-q substitute 10(8-q)+25q=155 80+15q=155 15q=75 q=5 d=3